OpenStudy (superhelp101):

Which of the following is the solution to the equation 25^2c = square root of 5(4c + 14)

OpenStudy (superhelp101):

let me rewrite it...

OpenStudy (superhelp101):

\[25^{2c}=\sqrt{5}^{4c+14}\]

OpenStudy (superhelp101):

@jim_thompson5910

OpenStudy (nick88888888):

Errrrr, here 25^(2c) = (√5)^(4c - 14) ((√5)^4)^(2c) = (√5)^(4c - 14) (√5)^(8c) = (√5)^(4c - 14) 8c = 4c - 14 4c = -14 c = -14/4 = Answer

OpenStudy (superhelp101):

that is 4c-14...I need 4c+14 ;)

OpenStudy (nick88888888):

Oh Mai Bad bru wait a sec

OpenStudy (superhelp101):

oh it's fine..never mind. it's like 2:40 here. I'm just gonna sleep ;) thx for the try thou

OpenStudy (superhelp101):

2:40 am

OpenStudy (nick88888888):

Oh its just 7/2 then, the answer would have been -7/2 but derp

OpenStudy (nick88888888):

Since i did 4c-14 not 4c+14 my bad for mis reading earlier, but yea its 7/2 then not -7/2

OpenStudy (superhelp101):

Lol didn't want to hurt any feelings ;)))

OpenStudy (nick88888888):

@superhelp101 ITs ok i dont get hurt easily ;3

OpenStudy (superhelp101):

Lol! ;)

OpenStudy (superhelp101):

So you live in S.K

OpenStudy (nick88888888):

Yea south korea :P

OpenStudy (superhelp101):

cool, gonna sleep now, k, thx, bye...

OpenStudy (nick88888888):

Bu bai, have a good sleep

OpenStudy (superhelp101):

okay thx! me best friend is from S.K too. :)